OpenStudy (anonymous):

Multiply and simplify: 2 radical 5 times the quantity 3 radical 2+3 radical 5. A.15 square root 10+6 square root 5 B.6 square root 10+6 square root 5 C.6 square root 10 +30 how do you solve this step by step ?

jhonyy9 (jhonyy9):

2sqrt5(3sqrt2+3sqrt5)=6sqrt10 +6sqrt(5*5) = 6sqrt10 +6*5 = 6sqrt10 +30 C. is right